OpenStudy (anonymous):

Show the expanded formx^4/x^2

OpenStudy (anonymous):

x^4/x^2 = x^2

OpenStudy (anonymous):

a^m/a^n = a^(m-n)

OpenStudy (anonymous):

x^4/x^2 = x^(4 - 2) = x^2 x^4/x^2 = x*x*x*x / x*x = x*x (x)(x)(x)(x) / (x)(x) = (x)(x) [x][x][x][x] / [x][x] = [x][x] {x}{x}{x}{x} / {x}{x} = {x}{x}